OpenStudy (anonymous):

v^2+64-16v= (v-8)(v+8)...right? I'm sorry for my neediness, I just have no one home to double check my work b4 submission! ;0

OpenStudy (saifoo.khan):

Just a bit mistake in signs

jimthompson5910 (jim_thompson5910):

no, you're thinking of the difference of squares formula

OpenStudy (anonymous):

(v-8)(v-8)

jimthompson5910 (jim_thompson5910):

or \[v^2+64-16v=(v-8)^2\] this is a perfect square

jimthompson5910 (jim_thompson5910):

so use the perfect square formula \[a^2-2ab+b^2=(a-b)^2\]

OpenStudy (saifoo.khan):

agree.

jimthompson5910 (jim_thompson5910):

in this case, a=v and b=8

OpenStudy (anonymous):

eiyiyi...yes!!